# Pursuing Top Growth with Novel Loss Function

### v1

@@ Abstract
Stock Selection , Time-Series Analysis , Convolutional Neural Network , Mixture of Experts , Sharpe Ratio , Cross-Entropy
[…]
@@ 1 Introduction
Stock price and movement prediction have always been extraordinarily challenging yet heavily sought-after tasks. Before the popularity of artificial intelligence and availability of unforeseen computing power present today, initial stages of our financial understanding consist of the Capital Asset Pricing Model (CAPM) (Sharpe,, 1964), the Efficient Market Hypothesis (EMH) (Fama,, 1970), and more. Decades of research following them have witnessed a vast number of articles that build upon these very fundamental concepts, including the 3, 4, and 5-factor models (Fama and French,, 1993; Carhart,, 1997; Fama and French,, 2015). In recent years, these theories still form the foundation underlying a significant amount of work such as the Markov Decision Process (Park et al.,, 2024) and ARIMA (Box et al.,, 2015), which have adopted more advanced mathematical and statistical techniques.
@@
The two main branches of techniques used to analyze stock prices in order to develop profitable trading strategies are fundamental and technical. The fundamental approach considers both internal factors such as earnings per share (EPS) in financial statements and external factors such as the macroeconomic environment and geopolitical conditions. Analyzing textual data such as news and online posting platforms in the field of Natural Language Processing (NLP) has also grown significantly in popularity in light of the success of Large Language Models (LLMs). There are many models that utilize word embeddings such as GloVe (Pennington et al.,, 2014; Zhang et al.,, 2022), BERT (Devlin,, 2018) or more (Lin et al.,, 2022) to process texts, which can be used as one component of the inputs or for sentiment analysis. The embeddings from LLMs can replace these traditional embedding techniques due to LLM’s much stronger language ability (Kim and Nikolaev,, 2024). On the other hand, our work falls under the technical side, where we primarily focus on historical price and volume data and the additional features that are constructed from them. To name a few examples utilizing primarily technical indicators, Hoseinzade and Haratizadeh, (2019) incorporates information from related markets as the third dimension of the data; Gezici and Sefer, (2024) converts time-series Exchange-Traded Funds (ETFs) data containing technical indicators into 2-dimensional images and uses vision transformers, image transformers, and Swin to make trading decisions. There are also many successful studies that use a combination of sentiment and technical analysis (Picasso et al.,, 2019; Prachyachuwong and Vateekul,, 2021).
@@
Many technical factors have been found and used over the years, and the market can become increasingly more efficient with respect to them (Chen and Zimmermann,, 2021). As a result, the linear relationship between the realized return and these factors is largely priced in. However, the inherent non-linear relationship among the factors and market movement is not captured easily, so stock market predictions remain a challenging task. To this end, recent advances in deep learning have led to the adoption of models such as Convolutional Neural Networks (CNN), Long Short-Term Memory (LSTM), and attention networks to enhance prediction accuracy (Lu et al.,, 2020; Luo et al.,, 2024; Zhao and Yuan,, 2023). While these models have shown promising results, most studies focus primarily on minimizing prediction errors rather than optimizing financial performance (Oncharoen and Vateekul,, 2018). Traditional loss functions fail to account for the financial impact of prediction errors, treating all errors equally rather than weighting them based on potential trading risks and returns.
Expository works (Ashtiani and Raahemi,, 2023; Patel et al.,, 2023) indicate that the choices of loss functions and evaluation metrics in the current deep learning literature are overwhelmingly repetitive. They are either classification metrics such as accuracy, entropy, precision, recall, and F-scores or regression metrics such as mean squared error (MSE), mean absolute percent error (MAPE), and root mean square error (RMSE). They each have their advantages in specific tasks, but a potential drawback of them is that they treat each prediction equally and then take some form of average. Even when they do well across the entire dataset, they may not be adept at practical financial applications such as finding the small number of top growth opportunities.
To address this limitation, recent research has explored improvements to loss functions in stock prediction. Zhao et al., (2024) proposes a generalized loss function that dynamically adjusts the cost of prediction errors based on data difficulty, achieving higher returns despite similar accuracy levels. Combining the advantages of both classification and regression metrics, we propose a novel loss function with a weighting scheme that prioritizes the model’s attention on those with high growth potential and high risk for steep loss. Specifically, the novel loss function is
where CE is cross-entropy and the weight $r_{\text{cap}}$ is the capped version of daily percentage return introduced in Section 3.3.1. We implement the novel loss function in CNN-based models, where stocks with higher percentage gains and losses are penalized more heavily during training. This allows us to align loss optimization with actual financial impact, leading to improved decision-making for traders and investors. The proposed method is evaluated using real-world stock market data, comparing its performance against linear regression models and traditional loss functions.
The highlights and key contributions of the paper are as follows.
@@
    We propose a CNN-based efficient trading system equipped with a novel return-weighted loss function that is specifically adept at capturing top growth opportunities.
@@
The remainder of the paper is organized as follows. In Section 2, we briefly formulate the objective of the study and how we achieve it. Section 3 contains data preparation and processing details. The main model architecture including the novel loss function is described in Section 4. We showcase the model’s performance and simulation results in Section 5 before the concluding Section 6. The code used for this study is continuously updated on Github.
@@
2 Study Overview
Given the daily features of a stock related to trading volume, momentum, volatility, and trend, which are collected and computed at the end of the day $T$, the goal of the study is to calculate a score that ranks all stocks for day $T+1$. This score determines how favorable it is to purchase each stock at the market open of day $T+1$ and hold it until at least by the market open of day $T+2$. The score is calculated every day after the market closes to produce a daily trading strategy. In simulating the various strategies in the analysis section later, we demonstrate that the top few scores are particularly more impactful and are most able to drive top growth.
The input $X$ for each stock on day $T$ is the 2-dimensional matrix
@@ 2 Study Overview
of size $m\times n$, where $n$ is the number of features and $m$ represents the number of days to which the value of the feature goes back. Namely, we consider the time series of each feature from day $T-m+1$ to day $T$ when predicting the action for day $T+1$.
@@
The score that determines our investment strategy is calculated in two steps. The first step is a model that maps the input data $X$ to a probability distribution $y_{pred}$ of 5 labels: Strong Sell, Sell, Hold, Buy, and Strong Buy. This is done by linear regressions in Section 4.2.1 and CNN models in Section 4.2.2 after introducing the novel return-weighted loss function in Section 4.1. Then in the second step, the score is calculated as the dot product between the predicted distribution $y_{pred}$ and the vector $[-2,-1,0,1,2]$. Lastly, we rank each stock according to the score that they receive. More details on how this is implemented and how all the models perform are in Sections 4 and 5, respectively.
3 Data Management
@@
This section is dedicated to presenting the data pipeline in the study. We obtain our data from the publicly available Yahoo Finance database. Given that our current study only requires basic daily information including price (open, close, high, low), trading volume, and sector information, other open source databases can be used as well. All stocks we select are from major US stock exchanges, such as the New York Stock Exchange (NYSE) and Nasdaq. We discuss how we obtain and process data in Section 3.1, construct several technical indicators/features from the previous step using the Python Technical Analysis (TA) library (Padial,, 2018) in Section 3.2, and how data are standardized and organized for training in Section 3.3. This is also partly represented in the left half of Figure 1. The data we use in this study are publicly available at Mendeley Data.
3.1 Data Sourcing and Processing
@@
We download daily stock price (open, close, high, low) data, volume (number of shares traded) data, and sector information from Yahoo Finance using its Python API. There are 11 sectors in the dataset, and each stock’s sector information is stored as a one-hot vector in 12 dimensions, where the extra dimension is for those without sector information in Yahoo Finance (such as ETFs). Several additional features are constructed from the price and volume data in Section 3.2, and sector information is incorporated into the feature dimension through an embedding. This resembles how the positional embeddings in the transformer architecture are added to the input (Vaswani et al.,, 2017), but in our case, the same embedding independent of the point in time is added to the feature dimension. We intentionally choose two time periods 2006-2010 and 2019-2024 to include the 2008 financial crisis and its aftermath, the 2020 COVID pandemic, and the heightened fluctuations observed in the market starting around 2022 due to high inflation, fear of recession, and geopolitical factors.
After obtaining data from Yahoo Finance, we filter stocks using the following steps. We first remove stocks if their average daily dollar trading volume is below 10 million. This is to avoid practical issues where orders cannot be filled promptly due to low volume, as in the simulation, we assume we are able to buy and sell stocks as soon as the market opens. Certain stocks experience significant price and trading volume drops starting at some point in the selected time period. Their price can become very low and highly volatile, and it is not appropriate to assume that all orders can be executed promptly at the desired price, even though these stocks can still maintain an average dollar trading volume over 10 million across the selected time period. Therefore, to eliminate this potential source of inaccuracies when simulating returns for these stocks, we manually change the daily return of a stock to 0 when its price drops below \$0.1 per share, and we consider the stock dead once it happens. Dead stocks are still in the dataset, but after the delegation, they do not contribute to our proposed return-weighted loss function (see definition in Section 4.1) or the simulation results significantly (see details in Section 5). Note that the models may still recommend to buy these stocks, and their returns would be recorded as usual before they die and recorded as zero after. There are only 2 dead stocks from 2006-2010 and 6 from 2019-2024, so the change of their daily return to 0 does not negatively impact the quality of the data. In this way, we also keep the number of stocks identical every day in each of the two time periods. This process results in 924 stocks in the 2006-2010 period with 1360 trading days for backtest and 2152 stocks in the 2019-2024 period with 1340 trading days for backtest.
3.2 Feature Engineering

### v2

@@ Abstract
Stock Selection, Loss Function, Deep Learning, Time-Series Analysis, Embeddings, Sharpe Ratio, Cross-Entropy
@@ 1 Introduction
Stock price and movement prediction have always been extraordinarily challenging yet heavily sought-after tasks. Before the popularity of artificial intelligence and availability of unforeseen computing power present today, initial stages of our financial understanding consist of the Capital Asset Pricing Model (CAPM) [33], the Efficient Market Hypothesis (EMH) [8], and more. Decades of research following them have witnessed a vast number of articles that build upon these very fundamental concepts, including the 3, 4, and 5-factor models [9, 3, 10]. In recent years, these theories still form the foundation underlying a significant amount of work such as the Markov Decision Process [27] and ARIMA [2], which have adopted more advanced mathematical and statistical techniques.
@@
The two main branches of techniques used to analyze stock prices in order to develop profitable trading strategies are fundamental and technical. The fundamental approach considers both internal factors such as price to earnings (PE), earnings per share (EPS) in financial statements and external factors such as the macroeconomic environment and geopolitical conditions. Analyzing textual data such as news and online posting platforms in the field of Natural Language Processing (NLP) has also grown significantly in popularity in light of the success of Large Language Models (LLMs). There are many models that utilize word embeddings such as GloVe [29, 42], BERT [5] or more [22] to process texts, which can be used as one component of the inputs or for sentiment analysis. The embeddings from LLMs can replace these traditional embedding techniques due to LLM’s much stronger language ability [18]. On the other hand, our work falls under the technical side, where we primarily focus on historical price and volume data and the additional features that are constructed from them. To name a few examples utilizing primarily technical indicators, Hoseinzade and Haratizadeh [13] incorporated information from related markets as the third dimension of the data; Gezici and Sefer [12] converted time-series Exchange-Traded Funds (ETFs) data containing technical indicators into 2-dimensional images and uses vision transformers, image transformers, and Swin to make trading decisions. There were also many successful studies that use a combination of sentiment and technical analysis [28, 30].
@@
Many technical factors have been found and used over the years, and the market can become increasingly more efficient with respect to them [4]. As a result, the linear relationship between the realized return and these factors is largely priced in. However, the inherent non-linear relationship among the factors and market movement is not captured easily, so stock market predictions remain a challenging task. To this end, recent advances in deep learning have led to the adoption of models such as Convolutional Neural Networks (CNN), Long Short-Term Memory (LSTM), and attention networks to enhance prediction accuracy [21, 23, 43]. While these models have shown promising results, most studies focus primarily on minimizing prediction errors rather than optimizing financial performance [24]. Traditional loss functions fail to account for the financial impact of prediction errors, treating all errors equally rather than weighting them based on potential trading risks and returns.
To address this limitation, we propose a novel enhancement to loss functions in the context of non-linear machine learning (ML) models. By incorporating return-weighted loss functions, where prediction errors for stocks with higher absolute returns are penalized more heavily, the model aims to align loss optimization with actual financial impact, especially those with high growth potentials. The highlights and key contributions of the paper are as follows.
@@
    We propose a ML-based efficient trading system equipped with a novel return-weighted loss function that is specifically adept at capturing top growth opportunities.
@@
The remainder of the paper is organized as follows. We mention related works in Section 2. In Section 3.1, we briefly formulate the objective of the study and the setup we establish to achieve it. Section 4 contains data, model training details, and various analysis to demonstrate the superiority of our proposed novel loss function. We conclude the study with a brief discussion of its limitations and provide several future direction in Section 5.
@@
2 Related Works
There have been undoubtedly a tremendous amount of literature studying the application of deep learning in financial markets. Early works such as [7] tested CNNs on stock direction forecasting, which achieved accuracies of 60–65%, although the result did not clearly surpass traditional methods like logistic regression or multilayer perceptrons. Hoseinzade and Haratizadeh [13] demonstrated that CNNs can help obtain useful features from diverse market variables which were then used to improve high-frequency forecasting on the CSI 300 index, especially under extreme market conditions such as the two time periods we choose in this study as well. Machine learning methods can supplement each other and benefit from other theoretical concepts. For instance, Durairaj and Mohan [6] combined chaos theory with CNNs to enhance predictive accuracy. Trivedi and Sangeetha [36] converted basic price data (open, close, high, low) into inputs compatible with pre-trained language models so that language models are able to process both textual and numerical information. They later fine-tuned language models to make stock trend predictions. Zeng et al. [40] introduced a CNN–Transformer hybrid that outperformed ARIMA and DeepAR on intraday S&P 500 data, using CNNs on short-term features and transformers to model long-term dependencies.
In addition to directional forecasting, other studies focused on linking prediction to trading performance. Jiang and Liang [16] applied CNNs within a reinforcement learning framework for cryptocurrency portfolio management, directly outputting portfolio weights to maximize returns, but they noted that the Sharpe ratios are not stable. More recently, Wang et al. [38] proposed a self-adaptive trading strategy that dynamically switches forecasting models and trading types, achieving both higher returns and lower drawdowns in the Hong Kong equity market.
We observe that, however, detailed studies on the loss function that models optimize are largely absent from existing work. Comprehensive surveys [35, 39] pointed out the common loss functions are regression metrics such as mean squared error (MSE) and mean absolute percent error (MAPE) and classification metrics such as cross entropy (CE). Each of these aforementioned loss functions have a vast number of possible variations adapting to specific tasks. Our proposed loss function is like a weighted cross entropy loss, with the weight on the granular level of each data point. However, typical weighted loss functions are typically used on the whole classes of labels to alleviate class-imbalance issues [20].
In the financial domain specifically, expository works [1, 26] on the role of ML in finance indicate that the choices of loss functions (and evaluation metrics) in the current deep learning literature are indeed overwhelmingly repetitive. They each have their advantages in specific tasks, but a potential drawback of them is that they treat each prediction equally and then take some form of average. Even when they do well across the entire dataset, they may not be adept at practical financial applications such as finding the small number of top growth opportunities. Some recent research has explored improvements to loss functions in stock prediction. A generalized loss function is proposed that dynamically adjusts the cost of prediction errors based on data difficulty, achieving higher returns despite similar accuracy levels [41]. Combining the advantages of both classification and regression metrics, we propose a novel loss function with a weighting scheme that prioritizes the model’s attention on those with high growth potential and high risk for steep loss. Specifically, the novel loss function is
where CE is cross-entropy and the weight $r_{\text{cap}}$ is the capped version of daily percentage return introduced in Section 3.1.1. We implement the novel loss function in ML-based models, where stocks with higher percentage gains and losses are penalized more heavily during training. This allows us to align loss optimization with actual financial impact, leading to improved decision-making for traders and investors. The proposed method is evaluated using real-world stock market data, comparing its performance against linear regression models and traditional loss functions.
3 Study Design based on Novel Loss Function
3.1 Problem Formulation and Strategy Overview
We propose a novel loss function and use it to develop trading systems that produce a score for every stock considered on a daily basis. These scores are calculated after the end of a trading day $T$ to rank all stocks and recommend the top stocks to invest in on the next trading day $T+1$. To calculate the score for a stock, we input into the model the following 2-dimensional matrix
@@ 3.1 Problem Formulation and Strategy Overview
of size $m\times n$, where $n$ is the number of technical features we consider, and $m$ represents the number of days to which the value of the feature goes back. Namely, we consider the time series of each feature from day Tm+1T-m+1 to day $T$ when predicting the action for day $T+1$.
The main component of our trading systems is a neural network consisting of various structures taking technical features as inputs. There are five initial price and volume features that are easily and publicly accessible (open, close, high, low, trading volume). We compute 27 additional features related to trading volume, momentum, volatility, and trend that are computed from the five initial features (See Table LABEL:tab:all_features for some of their description in the Appendix). Disregarding the four initial price features since they are unlikely closely correlated with stock return, we obtain $n=28$ features as the input of the model.
@@
The neural networks are the first step in calculating the score that determines our investment strategy. Specifically, we train neural networks guided by a novel loss function to map the input data $X$ to a probability distribution $y_{pred}$ of 5 labels: Strong Sell, Sell, Hold, Buy, and Strong Buy, respectively. In the second step, the score is calculated as the dot product between the predicted distribution $y_{pred}$ and the vector [2,1,0,1,2][-2,-1,0,1,2]. We rank each stock after combining the scores from three independent models and invest in the highest ranked stocks equally.
3.1.1 First Step of Loss Function - Daily Return
The practical implementation of our novel return-weighted loss function begins with the calculation of daily returns. The daily return $r_{d,T}$ (or $r_{d}$ if the day $T$ is implicit) of a stock is the percentage change from the opening price $p_{open,T+1}$ of day $T+1$ to the opening price $p_{open,T+2}$ of day $T+2$. We calculate $r_{d,T}$ as
$r_{d,T}=\frac{p_{open,T+2}-p_{open,T+1}}{p_{open,T+1}},$ (1)
which is then used to assign the following labels in the data as follows
  • $r_{d,T}\geq 3\%$ - Strong Buy - $y_{true}=[0,0,0,0,1]$,
  • $1\%<r_{d,T}<3\%$ - Buy - $y_{true}=[0,0,0,1,0]$,
  • $1\%<r_{d,T}\leq 1\% -$ Hold - $y_{true}=[0,0,1,0,0]$,
  • $3\%<r_{d,T}\leq-1\% -$ Sell - $y_{true}=[0,1,0,0,0]$,
  • $r_{d,T}\leq-3\%$ - Strong Sell - $y_{true}=[1,0,0,0,0]$.
The data distribution after the labeling is (7.43%, 18.50%, 46.73%, 19.45%, 7.88%) in the 2005-2010 period and (7.79%, 17.61%, 47.83%, 18.92%, 7.85%) in the 2019-2024 period. It signals a slightly positive skew in the data which can also skew the model predictions to be slightly more positive, which is discussed later in Section 4.3. The daily return $r_{d,T}$ is the objective for the three baseline linear regression models and another baseline CNN with MSE loss introduced in Section 3.4. In the formulation of the new loss function, $r_{d,T}$ serves as the weight for the cross-entropy loss that we describe in Section 3.2.
3.2 Novel Loss Function
Both the data and the loss function play a significant role in ML model’s performance. While data provide the resources for model training, the loss function that an ML model minimizes is a guide that directly impacts the performance and (downstream) applications of the model. Typical loss functions include Mean Squared Error (MSE) or its variants in the case of asset pricing and cross-entropy (CE) in the case of classifications. There are three notable considerations in using the very widely used MSE loss in the current study as follows.
  1. 1.
  2. 2.
  3. 3.
The other common loss function cross-entropy [32] is a measure of the difference between two probability distributions, so it is excellent at classification tasks. The cross-entropy of a predicted distribution $q=(q_{1},\dots,q_{n})$ relative to the true distribution $p=(p_{1},\dots,p_{n})$ is
$\mathrm{CE}(p,q)=-\sum_{i=1}^{n}p_{i}\log q_{i},$ (2)
The cross-entropy loss can treat incorrect classifications equally even when they have vastly different financial implications. For example, mislabeling Strong Sell and mislabeling Hold could lead to the same cross-entropy loss. Suppose there are two true labels Strong Sell (next day return $r_{1}=-5\%$, label $y_{1}=[1,0,0,0,0]$) and Hold (next day return $r_{2}=0.5\%$, label $y_{2}=[0,0,1,0,0]$), and that there are two predictions $\hat{y_{1}}=[0.2,0.1,0.1,0.1,0.5]$ and $\hat{y_{2}}=[0.1,0.1,0.2,0.5,0.1]$. By the definition of cross-entropy (2), $\text{CE}(y_{1},\hat{y_{1}})=\text{CE}(y_{2},\hat{y_{2}})=-\log(0.2)$, but in practice, these two mistakes are not considered the same in terms of their financial consequence. Therefore, we propose a weighting scheme for the cross-entropy loss function that naturally prioritizes stocks with large price movement so that the model will learn to focus on predicting those stocks correctly. To protect the models from being overly skewed by large values of daily percentage return $|r_{d}|$, we use a capped version of the daily return $r_{d}$ as the weight. The definition of the capped version $r_{\text{cap}}$ with cutoff at $50\%$ is
Our novel return-weighted loss function is
where $y_{\text{true}}$ and $y_{\text{pred}}$ are probability distributions over five labels.
Note that in the example above, our new loss has that $\text{loss}(y_{1},\hat{y_{1}})=-0.05\log(0.2)$, which is 10 times as much as $\text{loss}(y_{2},\hat{y_{2}})=-0.005\log(0.2)$. This aligns with our understanding that incorrectly classifying Strong Sell stocks is worse than incorrectly classifying Hold stocks. In fact, since the stocks labeled as Strong Sell or Strong Buy in the dataset both have their absolute daily percentage change greater than 3%, they always contribute to the loss function more significantly than those with the other three labels. Therefore, by minimizing loss during training, the models will identify and prioritize stocks with immediate breakout and breakdown potentials. Most stocks do not move significantly on a daily basis, and they do not increase the loss as much even when they are classified incorrectly. This leads to better detection for top growth opportunities but could lead to slightly worse overall ranking ability which we discuss in Section 4.3.
3.3 Model Architecture
We describe various neural network setups in this section and when applicable, the specific configurations that we use in this study.
Multi-Layered Perceptrons (MLPs): Also called Deep Neural Networks (DNNs), MLPs are relatively simple supervised machine learning algorithms in terms of data preparation and implementation. Given the versatility of these neural networks in general, they have been deployed in lots of financial literature and have achieved great success [11]. On the other hand, one key drawback of DNNs is that they only take 1-dimensional inputs. While it is possible to collapse multi-dimensional data into 1 dimension, DNNs do not have an inherent understanding of higher dimensional data structure. During learning, they might be able to pick up on the dimensionality in later hidden layers, but this is far from guaranteed and is at the cost of resources.
Convolutional Neural Networks (CNNs): CNNs are originally developed for tasks involving images, such as classification [19], segmentation [31] and more. Colored images are naturally 3-dimensional data: their three RGB channels being one dimension, and the height and width being the other two dimensions. For each color channel in the input data, convolutional layers apply a sliding window of fixed size $l$ by $w$ (also called kernel) and compute a linear combination of the data values in each window. The coefficients (also called weights) used in the combinations are trainable and are shared across all kernel positions within each input channel. Convolutions use separate weights for different input channels. One output channel is obtained by summing the results across input channels at each position. If there are multiple output channels, this process is repeated with different sets of weights for each output channel. One premise of using convolutions is that data points that are in the same kernel interact with each other in some way. In our study, we perform 1D convolutions on the time dimension, so the kernel has one dimension of size $l$, and this setup helps the model find patterns in each feature’s trend over time. Since there may not be a reasonable way to order the features such that all nearby features interact with each other, we do not convolve the features.
Attention: Introduced by Vaswani et al. (2017), the attention mechanism excels at language related and generative tasks. To name a few popular applications, attention is a crucial component in large language models and diffusion models. Given a sequential input such as a sentence or a time series of stock data in our case, we first construct contextual questions (Q) and keys (K) as attempted answers. The interactions between the questions and keys weight the values (V) to produce the output of one attention head. Multiple independent attention heads can be constructed in parallel forming Multi-Head Attentions (MHA), and their outputs are concatenated before taken to subsequent dense layers. The MHA layer and its subsequent layers make up one encoder block. Residual connections [14] are used among layers in the encoder block to speed up data passage. There are typically multiple consecutive encoder blocks in practice.
Sector Embedding: We embed a trainable one-hot vector calculated from the 12 sector dimensions into $n=28$ dimensions. The embedding is added to the third (feature) dimension of the input data $X$ of shape (None, $m$, $n$), and the embedding matrix is trained together with all other parameters. The way in which the sector embedding is added to the input data resembles how positional embeddings are used in transformer models [37]. In this particular application in stock selection, there are several reasons for why we need to use sector information to guide the technical features. The features we constructed for each stock should be interpreted differently in different sectors. Certain features have varying levels of impact depending on the sector they are in, and some features exhibit different levels of longevity and periodicity. Moreover, the embedding matrix only contributes $12\times 28=336$ additional parameters, serving as an economical method to integrate categorical data with continuous data.
For non-ML models, no sector embedding is added. For CNN models, after the sector embedding is added to the input data, the result goes through several layers of 1D convolutions without padding on the time dimension, thereby reducing the time dimension and increasing the feature dimension of the data. We eventually remove the time dimension through a global average pooling layer before passing the result to a series of dense layers. For attention-based models, the input data first go through three encoder blocks whose output has the same dimension as the input. Then, sector embedding is added as in the CNN models and the subsequent structures are the same. See Figure 2 for an illustration.
Figure 2: CNN and attention model architecture equipped with novel loss function
3.4 Baseline Methods
To obtain baseline performance for comparison, we choose several sets of features to conduct linear regression on. Since linear regression methods do not directly fit to the same probability distribution objective of the ML models, the objective of linear regression is the daily return $r_{d}$ (1). The sets of chosen features that we perform regression on are
  1. 1.
    3-day momentum, number of shares traded;
  2. 2.
    3-day momentum, relative strength index;
  3. 3.
    3-day momentum, 50-day moving average;
  4. 4.
    2-day momentum, 5-day momentum;
  5. 5.
    a subset of 12 non-technical features;
  6. 6.
    all 28 features.
We perform ordinary least squares (OLS), ridge and lasso with cross validation on these six sets of features. The data used for linear regression are identical to those for the ML models.
4 Implementation, Investment Simulation and Analysis
4.1 Data Management
@@
This subsection is dedicated to presenting the data pipeline in the study. We obtain our data from the publicly available Yahoo Finance database. Given that our current study only requires basic daily information including price (open, close, high, low), trading volume, and sector information, other open source databases can be used as well. All stocks in the dataset are from major US stock exchanges, such as S&P 500, New York Stock Exchange (NYSE), and Nasdaq. The data we use in this study are publicly available at Mendeley Data.
4.1.1 Data Sourcing and Processing
@@
We download daily stock price (open, close, high, low) data, volume (number of shares traded) data, and sector information from Yahoo Finance using its Python API. There are 11 sectors in the dataset, and each stock’s sector information is stored as a one-hot vector in 12 dimensions, where the extra dimension is for those without sector information in Yahoo Finance (such as ETFs). We intentionally choose two time periods 2006-2010 and 2019-2024 to include the 2008 financial crisis and its aftermath, the 2020 COVID pandemic, and the heightened fluctuations observed in the market starting around 2022 due to high inflation, fear of recession, and geopolitical factors. We process stocks by removing microcap stocks and those with low dollar trading volume. This results in 924 stocks in the 2006-2010 period with 1360 trading days for backtest and 2152 stocks in the 2019-2024 period with 1340 trading days for backtest.
4.1.2 Feature Engineering